# Generalizing multiple memories from a single drive: The hysteron latch

### v3

@@ Abstract
The single-dial combination lock is a mechanism for storing multiple values from a single input. By alternating between clockwise and counterclockwise rotation, the operator encodes the combination values as a series of turning points. Each new turning point must be nested within the previous two, so that the lock verifies not only the values but their exact sequence. Information about the stored values can be recovered by observing the resistance to further rotation  (?). Finally, erasure can be achieved with a large twist of the dial. Every operation is accomplished with a single control.
This elegant idea was known by 1909  (?), but it was rediscovered decades later as return-point memory, a generic behavior of many materials and systems that contain hysteretic elements—even those that appear quite different from the series of wheels in a combination lock  (?, ?, ?, ?). For example, an amorphous solid that is repeatedly sheared back and forth contains many localized groups of rearranging particles  (?, ?, ?). In experiments, forward and reverse shear drive each of these groups to switch its configuration reversibly between a “forward” (“$+$”) and “reverse” (“--”) state—but only when a sufficiently large deformation is applied in each direction, meaning that the rearrangements have hysteresis. Collectively, the states of these rearrangements uniquely encode the nested turning points of shear strain (illustrated in Fig. 1A), and a suitable readout protocol can recover some or all of that history  (?, ?). Unlike in the combination lock, the memory is explained by the observation that rearranging groups act as “hysterons,” bistable elements that are the building blocks of many models of hysteresis  (?). The way return-point memory arises in ensembles of independent hysterons is rigorously understood  (?, ?), and the same picture emerges from observations of creases in crumpled sheets  (?), buckling units in mechanical metamaterials  (?, ?), and from models of rocks  (?), ferromagnetic, ferroelectric and martensitic materials  (?), and many more systems  (?, ?, ?).
@@
In some of these systems, interactions between hysterons are also apparent. In the case of the amorphous solid, the minimum imposed strain needed to drive a group of rearranging particles into the “$+$” state might be raised or lowered according to the state of another nearby group, due to their coupling via the material between them. Of particular interest are “frustrated” or antiferromagnetic interactions that make it harder to drive a pair of hysterons into the same state. When present among large numbers of hysterons or other elements that relax, frustrated interactions are associated with the extraordinarily rugged energy landscapes of glassy materials like amorphous solids and crumpled sheets  (?, ?). Recent simulations of hysterons with frustrated interactions have explored intriguing behaviors that are inconsistent with return-point memory, such as counting cycles of driving  (?, ?, ?, ?), and interacting mechanical hysterons have now been demonstrated in experiments  (?, ?, ?, ?, ?).
Theoretically, the behaviors of interacting hysterons can be arbitrarily complex computations, given enough hysterons and the precise control of their parameters  (?, ?). Yet the example of return-point memory shows that a useful and capacious memory behavior—the ability to store sequences of turning points—can instead arise from an ensemble of simple units (hysterons), making it ubiquitous in both designed and disordered systems. Encoding longer sequences merely involves expanding the ensemble size, much as one would add more wheels to the mechanism of a combination lock. This example leads us to ask whether, among the plethora of behaviors enabled by hysterons’ interactions, there is an analogous form of memory based on repeating a minimal unit.
In this paper, we show that return-point memory has a direct analogue in hysterons with frustrated interactions. While return-point memory stores the turning points of symmetric driving (Fig. 1A), frustrated interactions let hysterons store the turning points of asymmetric or rectified driving (Fig. 1B), such as when a bridge is crossed by a series of vehicles with different weights. In contrast, any system with pure return-point memory, including the combination lock, can store only the largest turning point of asymmetric driving—a fact we verify by attempting to read out memories in simulations.
Remarkably, just as the hysteron is the repeated unit that gives rise to return-point memories, we find that the unit for nested memories of asymmetric driving is nearly as elemental: just two hysterons with a frustrated interaction. This pairwise mechanism was recently confirmed experimentally by Paulsen, with hysterons realized as two bistable rotors coupled via springs to each other and to a driving rod  (?). From this basic unit of memory, we formulate a general principle of “latching”: conditions for preventing the system from returning to its previous state, that allow multiple memories to be stored with either symmetric or asymmetric driving. Our results point the way to new studies of interacting relaxations in disordered matter, and to the rational design of new information-processing capabilities in mechanical systems.
Model — We are motivated by models of disordered solid materials and metamaterials, wherein a system contains mesoscopic “hysterons” that can reversibly switch between two states under an imposed strain $\gamma$—for example by rearrangements of particles in an amorphous solid  (?, ?, ?), or snap-through of a buckled beam  (?, ?). These models capture many mesoscopic details of experiments and simulations  (?, ?, ?, ?, ?, ?, ?), especially when hysterons can interact. The $i$th hysteron switches its state $S_{i}$ at thresholds $\gamma_{i}^{\pm}$: when $\gamma>\gamma_{i}^{+}$, $S_{i}=+1$; when $\gamma<\gamma_{i}^{-}$, $S_{i}=-1$; and when $\gamma_{i}^{-}<\gamma<\gamma_{i}^{+}$, $S_{i}$ remains in its previous state. For convenience we write states as “$+$” and “--”. In our simulations, the “system” is an ensemble of many groups of $N$ hysterons with random parameters, modeling interactions via perturbed thresholds
@@
where $S_{j\neq i}$ represents the states of all hysterons excluding hysteron $i$, and $J_{ij}$ is the $N\times N$ interaction matrix. This is equivalent to the model in Ref. ?. $J_{ij}<0$ represents a frustrated interaction in the sense that one hysteron flip inhibits another hysteron from flipping in the same direction. Large, disordered systems with these interactions tend to be “glassy,” in that they have rugged energy landscapes with many metastable states that are far from a global energy minimum  (?). We initially omit positive (cooperative) interactions that cause one hysteron’s flip to encourage like flips and that can lead to avalanches, since the memory behavior with those interactions is known to be qualitatively the same as with non-interacting hysterons  (?). Accordingly, we choose $J_{ij}$ with uniform probability from $[-J_{0},0]$, with $J_{0}=0.01$ unless otherwise specified. Each hysteron’s $\gamma^{\pm}_{i}$ are chosen by drawing two values from a uniform distribution on $[-0.1,0.1]$, and then ordering them so that $\gamma_{i}^{-}$ $<$ $\gamma_{i}^{+}$, corresponding to rearrangements that dissipate energy.
@@
To simulate this model as $\gamma$ is varied, we use the open-source hysteron software package  (?). Like molecular dynamics simulations of amorphous solids, we work in the athermal and quasistatic limit: the algorithm identifies the hysteron that will flip soonest as $\gamma$ changes, and then holds $\gamma$ fixed while it updates any other hysterons that were destabilized via interactions (e.g. an avalanche), starting with the hysteron farthest past its threshold. When $N>2$, the random parameters would occasionally prevent the algorithm from finding a stable state; these cases were discarded  (?). Unless specifically noted, we avoid the case where $J_{ij}$ and $J_{ji}$ have opposite signs, for which this issue is common.
1 Results
1.1 Memories in Non-Interacting and Interacting Systems
Protocols for initializing each simulation and writing memories are shown in Fig. 1; example readout protocols are in Fig. 2(A, C). We begin each simulation with all hysterons in the “--” state, as though $\gamma\to-\infty$, and then we drive the hysterons to $\gamma=0$ before encoding memories. (Starting at $\gamma\to\infty$ similarly supports our conclusions; it is considered in Fig. S1.)
@@ 1.1 Memories in Non-Interacting and Interacting Systems
Without loss of generality, in single-memory tests we take the stored strain amplitude to be 4%; in two-memory tests we use 4% and 2%; and in tests with asymmetric driving we take the bottom turning point to be 0. The amplitudes correspond to commonly-used strains in studies of memory in amorphous solids  (?, ?, ?, ?), but they can be rescaled arbitrarily to match other physical systems (Supplementary Text). After the writing cycle(s), the state is saved as $S_{\text{mem}}$. To read out the memories encoded in $S_{\text{mem}}$, we apply a series of cycles with increasing amplitude $\gamma_{\text{read}}=0,\ 0.005,\ 0.01\dots$. This is a “serial” protocol that is suited to experiments, as opposed to a “parallel” protocol in which a separate copy of the system is made for each readout cycle  (?). After each cycle with amplitude $\gamma_{\text{read}}$, we record the fraction of hysterons that are different from $S_{\text{mem}}$. We average this fraction over the entire ensemble and plot it against $\gamma_{\text{read}}$.
@@
Figure 3A shows readouts after the symmetric two-amplitude protocol labeled “4, 2” in Fig. 1A. The curve from an ensemble of non-interacting hysterons ($J_{ij}=0$) is reproduced from the “4, 2” curve of Fig. 2B; it features a local minimum that indicates the most recent memory ($2\%$) and a cusp that indicates the larger memory ($4\%$), just as in the corresponding amorphous solid experiments and molecular dynamics simulations  (?, ?, ?, ?). However, in Fig. 3B driving the same ensemble asymmetrically changes its behavior dramatically: the curve shows no signature of the smaller driving amplitude, rising rapidly only after $4\%$.
@@
These strikingly different behaviors are both expected from return-point memory, wherein the system remembers the turning points of driving. As long as $\gamma$ is bounded between any pair of turning points, visiting either turning point will return the system to the state it had when it was at that turning point before  (?, ?). This property is recursive, meaning that one may encode more than one memory with a symmetric driving protocol by decreasing the strain amplitude, such that each new pair of turning points is nested within the last as in Fig. 2A.
@@
We now consider frustrated interactions, $J_{ij}<0$, as found in models of “glassy” matter such as crumpled sheets, disordered or amorphous solids, or spin ice and spin glass. Frustration means that one relaxation inhibits others, leading to these materials’ characteristically rugged landscapes of metastable states with broad distributions of energy barriers  (?). In hysteron models, the sequences in which hysterons switch during forward or reverse shear become mutable, so that return-point memory is no longer assured  (?, ?, ?, ?). The red curves of Fig. 3A show that nonetheless, replacing the ensemble of single hysterons with an ensemble of frustrated pairs merely perturbs the return-point memory of symmetric driving. However, in Fig. 3B where the readout curve for asymmetric driving had been zero, there is now a clear signature of both memories. This new signal resembles the much larger one from symmetric driving, suggesting a connection.
1.2 Memory Mechanisms
@@
Figure 4: Mechanisms for memory. (A) Readouts of single memory of symmetric driving. Local minima indicate the memory. (B) Schematic of how a single hysteron contributes to memory, showing hysteron thresholds on the left and an abridged driving protocol on the right. Changes in state during readout are marked with circled $+$ and --. State at the end of each cycle is shown below; only the $+$ state contributes to readout. (C) Readouts of single memory with asymmetric driving. The curve without interactions is zero for $\gamma_{\text{read}}\leq 4\%$ and then increases, while the curve with frustration is non-monotonic. (D) Schematic for a frustrated pair with asymmetric driving. For clarity, we show only how the interaction splits the hysteron bottoms $\gamma^{-}_{1}$ and $\gamma^{-}_{2}$ as a function of the other hysteron’s state, since the ordering of these four values is crucial for non-zero readout. The hysteron tops $\gamma_{1}^{+}$ and $\gamma_{2}^{+}$ can vary widely (Supplementary Text).
Figure 5: Latching pair transition graphs and scaling. (A) Transition graphs that can give rise to non-zero readout below the remembered strain in simulations of frustrated hysterons. Transitions while increasing (decreasing) strain are in red (blue). Left graph corresponds to Fig. 4D. Thick arrow represents an avalanche in which ++-+ - transitions to +-+- + upon increasing strain, by way of the unstable intermediate state +++++ +. (B) Peak signal in Fig. 4C for $\gamma_{\text{read}}<4\%$, varying interaction scale $J_{0}$ and number of mutually-interacting hysterons $N$, scaled by the number of pair interactions per hysteron ($N-1$). In “$9\pm$,” half of the $J_{ij},J_{ji}$ pairs are positive (cooperative), so data are scaled by $(N-1)/2$. Each point is the average of $10^{9}N$ hysterons. Dashed line shows $J_{0}^{2}$ scaling for reference.
@@
Just as with two memories, in Fig. 4C frustration enables a single memory of asymmetric driving that resembles its symmetric counterpart. We show in Fig. 4D that this happens by an analogous mechanism involving a frustrated pair of hysterons. Frustration allows a two-hysteron latching behavior in which revisiting the turning point at $\gamma=0$ can fail to restore the previous state. The mechanisms of Figs. 4B and 4C are thus equivalent when one treats each whole cycle as one transition, either to the same state or to a new state  (?).
1.3 Scaling the Two-Hysteron Latch
@@
as in Fig. 4D. These inequalities were verified in simulations by testing $10^{9}$ random pairs. In addition to the explicit flipping thresholds shown, it can be useful to visualize the order in which states are visited via a “transition graph.” The graphs that can lead to latching in our frustrated simulations—and hence to non-zero readout—are shown in Fig. 5A. If we let $J_{ij}$ and $J_{ji}$ have opposite signs, a third transition graph can contribute (see figure S3 for details).
@@
The interaction strength sets the threshold “window” size $\gamma^{-}_{1}(+)-\gamma^{-}_{1}(-)=-2J_{12}$, into which both 0 and the interval $[\gamma_{2}^{-}(-),\gamma_{2}^{-}(+)]$ must fall. For ensembles with uniformly-distributed parameters like those reported here, these two requirements make the probability $P$ for Eq. 2 second-order in the interaction strength, i.e. $P\sim J_{0}^{2}$. The signals in Fig. 4C are thus small compared to the result from return-point memory under symmetric driving, which is zeroth-order in the sense that it may be obtained with $J_{0}=0$. The $J_{0}^{2}$ scaling is confirmed for small interaction strength ($J_{0}\ll 4\%$) in Fig. 5B, where we measure the maximum height of the readout signal for $\gamma_{\text{read}}<4\%$ (see Supplementary Text for details). In a design context, our analysis means that greatest tolerance for manufacturing errors—and for variation in the bottom turning point of driving—corresponds to large $J_{12}$ and small $J_{21}$, while keeping $\gamma^{-}_{1}(+)$ below the smallest amplitude to be remembered. Crucially, hysteron pairs that fail to satisfy Eq. 2 do not corrupt an ensemble’s memory; they are simply absent from readout.
@@
The scaling estimate presented above is a departure from analyses of $P$ based solely on the ordering of thresholds, as considered by van Hecke  (?): here we included the turning point of asymmetric driving, which additionally isolates the $J_{0}^{2}$ behavior by cutting out the zeroth-order response.
@@
Remarkably, the two-hysteron latch is also how memories of asymmetric driving arise in larger groups of interacting hysterons, so that this motif may be observed in a bulk disordered material or metamaterial  (?, ?). Figure 5B shows nearly identical results for larger, mutually-interacting groups after dividing out the multiplicity of frustrated pairs. These results hold even when we randomly make half of the interactions cooperative (drawn from $[0,J_{0}]$)—strongly suggesting that Fig. 4D is the dominant mechanism despite many more possible behaviors  (?, ?, ?, ?). In the Supplementary Text table S1 we further show that the memory-forming portions of these larger groups tend to have the same kinematics and interaction strengths as in $N=2$. This remarkable conservation is possible because hysterons that do not contribute to asymmetric readout are largely following return-point memory, and so their states and transitions vary little from cycle to cycle.
1.4 Nesting Memories of Asymmetric Driving

### v4

@@ Abstract
  • Short title:
    Generalizing multiple memories from a single drive
  • Teaser:
    Adding frustrated interactions within a system with hysteresis reveals a general principle for storing multiple values from the history of a single variable, as in a combination lock.
I Introduction
The single-dial combination lock is a mechanism for storing multiple values from a single input. By alternating between clockwise and counterclockwise rotation, the operator stores the combination values as a series of turning points. Each new turning point must be nested within the previous two, so that the lock verifies not only the values but their exact sequence. Information about the stored values can be recovered by observing the resistance to further rotation [1]. Finally, erasure can be achieved with a large twist of the dial. Every operation is accomplished with a single control.
This elegant idea was known by 1909 [2], but it was rediscovered decades later as return-point memory: a general principle for systems of hysteretic elements driven by a field that alternately increases and decreases [3, 4, 1]. Applications have ranged from ferromagnets [3, 5] to rocks [6], and more recently, memories of deformation in amorphous solids [7, 8], crumpled sheets [9], and designed metamaterials that feature origami folds or buckling beams [10, 11]. In these newer examples, observations tied each system’s memory of mechanical strain to the hysteretic elements within it: particle rearrangements, creases, or buckling units, all modeled as bistable “hysterons.”
In this paper, we consider a simple change to the driving that leaves the combination lock or any other system with return-point memory unable to store multiple values. In conventional symmetric driving values are introduced by changing both the upper and lower turning points of motion (Fig. 1a). In asymmetric driving one turning point is kept fixed (Fig. 1b), such as when a bridge is crossed by a series of vehicles with different weights. This simple change allows no memories but the largest one, a fact we verify by writing and reading memories in ensembles of model hysterons [12, 13, 7] (Fig. 2). However, in place of return-point memory arising from single hysterons, for asymmetric driving we find an equivalent mechanism based on coupled pairs of hysterons, allowing a system to store multiple driving amplitudes and ensuring that it always remembers the most recent input. We then formulate a general “latching” principle for storing multiple memories sequentially from a single drive, whether symmetric or asymmetric. Our results point the way to new studies of glassy physics in disordered matter, and to the rational design of new information-processing capabilities in mechanical systems.
@@ I Introduction
II Results
We are motivated by models of disordered solid materials and metamaterials, wherein a system contains mesoscopic “hysterons” that can reversibly switch between two states under an imposed strain $\gamma$—for example by rearrangements of particles in an amorphous solid [14, 8, 15], or snap-through of a buckled beam [10, 16]. These models capture many mesoscopic details of experiments and simulations [8, 7, 17, 18, 11, 10, 19], especially when hysterons can interact. The $i$th hysteron switches its state $S_{i}$ at thresholds $\gamma_{i}^{\pm}$: when $\gamma>\gamma_{i}^{+}$, $S_{i}=+1$; when $\gamma<\gamma_{i}^{-}$, $S_{i}=-1$; and when $\gamma_{i}^{-}<\gamma<\gamma_{i}^{+}$, $S_{i}$ remains in its previous state. For convenience we write states as “$+$” and “--”. In our simulations, the “system” is an ensemble of many groups of $N$ hysterons with random parameters, modeling interactions via perturbed thresholds
@@ II Results
where $S_{j\neq i}$ represents the states of all hysterons excluding hysteron $i$, and $J_{ij}$ is the $N\times N$ interaction matrix. This is equivalent to the model in Ref. [20]. $J_{ij}<0$ represents a frustrated interaction, where one hysteron flip inhibits another from flipping. We choose $J_{ij}$ with uniform probability from $[-J_{0},0]$, with $J_{0}=0.01$ unless otherwise specified. Each hysteron’s $\gamma^{\pm}_{i}$ are chosen by drawing two values from a uniform distribution on $[-0.1,0.1]$, and then ordering them so that $\gamma_{i}^{-}$ $<$ $\gamma_{i}^{+}$, corresponding to rearrangements that dissipate energy.
II.1 Memories in Non-Interacting and Interacting Systems
@@
Protocols for initializing each simulation and writing memories are shown in Fig. 1; example readout protocols are in Fig. 2(a, c). We begin each simulation with all hysterons in the “--” state, as though $\gamma\to-\infty$, and then we drive the hysterons to $\gamma=0$ before encoding memories. (Starting at $\gamma\to\infty$ similarly supports our conclusions; it is considered in the Supplemental Materials [21].) Simulations used the open-source hysteron software package [20]. Like molecular dynamics simulations of amorphous solids, we work in the athermal and quasistatic limit: the algorithm identifies the hysteron that will flip soonest as $\gamma$ changes, and then holds $\gamma$ fixed while it updates any other hysterons that were destabilized via interactions (e.g. an avalanche), starting with the hysteron farthest past its threshold. When $N>2$, the random parameters would occasionally prevent the algorithm from finding a stable state; these cases were discarded [20]. We did not consider cases where $J_{ij}$ and $J_{ji}$ have opposite signs, for which this issue is common.
@@
Without loss of generality, we take the single stored strain amplitude to be 4%, and in two-memory tests we use 4% and 2%. After the writing cycle(s), the state is saved as $S_{\text{mem}}$. To read out the memories encoded in $S_{\text{mem}}$, we apply a series of cycles with increasing amplitude $\gamma_{\text{read}}=0,\ 0.005,\ 0.01\dots$. This is a “serial” protocol that is suited to experiments, as opposed to a “parallel” protocol in which a separate copy of the system is made for each readout cycle [13]. After each cycle with amplitude $\gamma_{\text{read}}$, we record the fraction of hysterons that are different from $S_{\text{mem}}$. We average this fraction over the entire ensemble and plot it against $\gamma_{\text{read}}$.
@@
Figure 3a shows readouts after the symmetric two-amplitude protocol labeled “4, 2” in Fig. 1a. The curve from an ensemble of non-interacting hysterons ($J_{ij}=0$) is reproduced from the “4, 2” curve of Fig. 2b; it features a local minimum that indicates the most recent memory ($2\%$) and a cusp that indicates the larger memory ($4\%$), just as in the corresponding amorphous solid experiments and molecular dynamics simulations [22, 13, 7, 23]. However, in Fig. 3b driving the same ensemble asymmetrically changes its behavior dramatically: the curve shows no signature of the smaller driving amplitude, rising rapidly only after $4\%$.
@@
These strikingly different behaviors are both expected from return-point memory, wherein the system remembers the turning points of driving. As long as $\gamma$ is bounded between any pair of turning points, visiting either turning point will return the system to the state it had when it was at that turning point before [3, 4]. This property is recursive, meaning that one may encode more than one memory with a symmetric driving protocol by decreasing the strain amplitude, such that each new pair of turning points is nested within the last as in Fig. 2a.
@@
We now consider frustrated interactions, $J_{ij}<0$, as found in models of “glassy” matter such as crumpled sheets, disordered or amorphous solids, or spin ice and spin glass. Frustration means that one relaxation inhibits others, leading to these materials’ characteristically rugged landscapes of metastable states with broad distributions of energy barriers [24]. In hysteron models, the sequences in which hysterons switch during forward or reverse shear become mutable, so that return-point memory is no longer assured [4, 25, 26, 18]. The red curves of Fig. 3a show that nonetheless, replacing the ensemble of single hysterons with an ensemble of frustrated pairs merely perturbs the return-point memory of symmetric driving. However, in Fig. 3b where the readout curve for asymmetric driving had been zero, there is now a clear signature of both memories. This new signal resembles the much larger one from symmetric driving, suggesting a connection.
Figure 4: Mechanisms for memory. (a) Readouts of single memory of symmetric driving. Local minima indicate the memory. (b) Schematic of how a single hysteron contributes to memory, showing hysteron thresholds on the left and an abridged driving protocol on the right. Changes in state during readout are marked with circled $+$ and --. State at the end of each cycle is shown below; only the $+$ state contributes to readout. (c) Readouts of single memory with asymmetric driving. The curve without interactions is zero for $\gamma_{\text{read}}\leq 4\%$ and then increases, while the curve with frustration is non-monotonic. (d) Schematic for a frustrated pair with asymmetric driving. For clarity, we show only how the interaction splits the hysteron bottoms $\gamma^{-}_{1}$ and $\gamma^{-}_{2}$ as a function of the other hysteron’s state, since the ordering of these four values is crucial for non-zero readout. The hysteron tops $\gamma_{1}^{+}$ and $\gamma_{2}^{+}$ can vary widely; see [21] for details. (e) Transition graphs that can give rise to non-zero readout below the remembered strain in simulations of frustrated hysterons. Transitions while increasing (decreasing) strain are in red (blue). Left graph corresponds to (d). Thicker arrow represents an avalanche in which the intermediate state +-+- + falls immediately into --- -. (f) Peak signal in (c) for $\gamma_{\text{read}}<4\%$, varying interaction scale $J_{0}$ and number of mutually-interacting hysterons $N$, scaled by the number of pair interactions per hysteron ($N-1$). In “$9\pm$,” half of the $J_{ij},J_{ji}$ pairs are positive (cooperative), so data are scaled by $(N-1)/2$. Each point is the average of $10^{9}N$ hysterons. Dashed line shows $J_{0}^{2}$ scaling for reference.
II.2 Memory Mechanisms
@@
Just as with two memories, in Fig. 4c frustration enables a single memory of asymmetric driving that resembles its symmetric counterpart. We show in Fig. 4d that this happens by an analogous mechanism involving a frustrated pair of hysterons. Frustration allows a two-hysteron latching behavior in which revisiting the turning point at $\gamma=0$ can fail to restore the previous state. The mechanisms of Figs. 4b and 4c are thus equivalent when one treats each whole cycle as one transition, either to the same state or to a new state [27].
II.3 Scaling the Two-Hysteron Latch
@@
as in Fig. 4d. These inequalities were verified in simulations by testing $10^{9}$ random pairs. In addition to the explicit flipping thresholds shown, it can be useful to visualize the order in which states are visited via a “transition graph.” The graphs that can lead to latching in our frustrated simulations—and hence to non-zero readout—are shown in Fig. 4e. If we let $J_{ij}$ and $J_{ji}$ have opposite signs, a third transition graph can contribute (see Supplementary Materials for details [21]).
@@
The interaction strength sets the threshold “window” size $\gamma^{-}_{1}(+)-\gamma^{-}_{1}(-)=-2J_{12}$, into which both 0 and the interval $[\gamma_{2}^{-}(-),\gamma_{2}^{-}(+)]$ must fall. For ensembles with uniformly-distributed parameters like those reported here, these two requirements make the probability $P$ for Eq. 2 second-order in the interaction strength, i.e. $P\sim J_{0}^{2}$. The signals in Fig. 4c are thus small compared to the result from return-point memory under symmetric driving, which is zeroth-order in the sense that it may be obtained with $J_{0}=0$. The $J_{0}^{2}$ scaling is confirmed for small interaction strength ($J_{0}\ll 4\%$) in Fig. 4f, where we measure the maximum height of the readout signal for $\gamma_{\text{read}}<4\%$ [21]. In a design context, our analysis means that greatest tolerance for manufacturing errors corresponds to large $J_{12}$ and small $J_{21}$, while keeping $\gamma^{-}_{1}(+)$ below the smallest amplitude to be remembered. Crucially, hysteron pairs that fail to satisfy Eq. 2 do not corrupt an ensemble’s memory; they are simply absent from readout.
@@
The scaling estimate presented above is a departure from analyses of $P$ based solely on the ordering of thresholds, as considered by van Hecke [26]: here we included the turning point of asymmetric driving, which additionally isolates the $J_{0}^{2}$ behavior by cutting out the zeroth-order response.
@@
Remarkably, the two-hysteron latch is also how memories of asymmetric driving arise in larger groups of interacting hysterons, so that this motif may be observed in a bulk disordered material or metamaterial [16, 28]. Figure 4f shows nearly identical results for larger, mutually-interacting groups after dividing out the multiplicity of frustrated pairs. These results hold even when we randomly make half of the interactions cooperative (drawn from $[0,J_{0}]$)—strongly suggesting that Fig. 4d is the dominant mechanism despite many more possible behaviors [26, 20, 29, 18]. In the Supplemental Materials [21] we further show that the memory-forming portions of these larger groups tend to have the same kinematics and interaction strengths as in $N=2$. This remarkable conservation is possible because hysterons that do not contribute to asymmetric readout are largely following return-point memory, and so their states and transitions vary little from cycle to cycle.
Figure 5: Graphical analysis of multiple memories. Each latching pair like Fig. 4d corresponds to a point on a plane, according to its upper thresholds. (a) An infinite ensemble of two-hysteron latches with continuously-distributed thresholds, all initialized to the --- - state. The thresholds are labeled $\gamma_{A}$, $\gamma_{B}$ for convenience. In accordance with Fig. 4d, $\gamma_{B}>\gamma_{A}>0$. (b) “Template” for how an asymmetric cycle with amplitude $\gamma_{0}$ changes the states of latches in the ensemble. (c) The ensemble in (a), after one cycle with amplitude 4% forms a memory. (d) The ensemble encodes two memories after cycles with 4% and 2% amplitude. (e) Readout of the ensemble in (d), reporting the fraction of hysterons in different states, in latches that conform to Fig. 4d and/or Fig. 4e. Curves show readout of all latches (“all”), and of ensembles generated separately to match each labeled region in (d).
II.4 Nesting Memories of Asymmetric Driving